# SRD: Reinforcement-Learned Semantic Perturbation for Backdoor Defense in VLMs

### v1

@@ Abstract
Vision-Language Models (VLMs) have made significant progress in the task of image caption generation, yet recent research has found that they are highly vulnerable to backdoor attacks. An attacker can activate anomalous behaviors within the model at test time by injecting imperceptible perturbations (e.g., local pixel triggers or global semantic phrases) into the training data, thereby causing the model to generate attacker-specified malicious descriptions for specific inputs. These attacks are particularly challenging to detect and defend against due to the stealthiness and cross-modal propagation of the trigger signals. In this paper, we identify two key vulnerabilities by analyzing existing attack samples: (1) the model exhibits abnormal attention concentration on certain regions of the input image, and (2) backdoor attacks often induce semantic drift and sentence incoherence. Based on these insights, we propose Semantic Reward Defense (SRD), a reinforcement learning framework that mitigates backdoor behavior without requiring any prior knowledge of trigger patterns. SRD learns to apply discrete perturbations (such as occlusion, color masking, and other localized modifications) to sensitive contextual regions of the image input via a Deep Q-Network strategy, aiming to divert attention and block the activation of malicious pathways. To guide policy optimization, we design a reward signal named semantic fidelity score, which jointly assesses the semantic consistency and linguistic fluency of the generated captions, encouraging the agent to achieve a robust yet faithful output. Experiments on mainstream VLMs across datasets demonstrate that SRD can significantly reduce the success rate of both local and global backdoor attacks while maintaining generation quality on clean inputs. The algorithm effectively reduces the attack success rate to 5.6%, with a performance drop of no more than 10% on clean inputs. We provide a trigger-agnostic and policy-interpretable paradigm for defending multimodal generative models against stealthy backdoor threats.
1 Introduction
Visual Language Models (VLMs) awadalla2023openflamingo ; zhu2023minigpt ; chung2024scaling have demonstrated strong cross-modal reasoning capabilities in tasks such as image-text matching and generation and have been widely adopted in real-world applications such as search and recommendation systems, computer-aided diagnosis, and multimodal question answering. Nevertheless, VLMs still face serious security challenges from backdoor attacks xu2024shadowcast ; lyu2024backdooring ; lyu2024trojvlm ; liang2025vl ; gu2019badnets ; liang2023badclip ; liang2024vl ; zhu2024breaking ; liang2024revisiting ; liu2024compromising ; liu2025natural . In such attacks, the attacker injects specific trigger patterns into images or text during the training phase so that the model outputs the attacker’s predetermined results once these patterns are detected during inference, e.g., misrepresenting an arbitrary image as “safe and harmless” xu2024shadowcast ; nguyen2020input ; tao2024imgtrojan ; liang2024revisiting .
@@ 1 Introduction
In recent years, backdoor attacks carlini2023aligned ; qi2024visual ; wang2024white ; niu2024jailbreaking ; gu2024agent ; lu2024test ; ni2024physical ; liu2024survey against VLMs have evolved, mainly including trigger-based attacks (embedding explicit or implicit patterns in images) and global perturbation-based attacks (injecting imperceptible noise to manipulate behavior). These attacks usually do not change the image appearance but can manipulate the model output at the semantic level, posing a serious challenge to traditional detection mechanisms. Despite the growing prominence of their potential threats, existing defense research wang2022universal ; liang2024unlearning ; kuang2024adversarial ; guo2024copyrightshield ; wang2025lie is still relatively limited against poisoned data.
However, our experimental analysis show that existing backdoor attacks still expose critical weaknesses at the cross-modal alignment mechanism and text output levels. First, we observe significant cross-modal anomalous attention coupling in the model when the backdoor is triggered. Visualization via attention heatmaps shows that the model’s attention tends to focus abnormally on the region in the image where the trigger is located and establishes a strong association with the target text. This anomalous coupling suggests that the triggers are able to manipulate the model’s attention at the visual level and guide it toward cross-modal alignment with the attacker’s predefined semantic targets, thus activating the backdoor. Second, the anomalous output of the model suffers from a lack of semantic fidelity in the text generation results, which is mainly manifested as semantic drift and semantic inconsistency. On the one hand, the generated text tends to revolve around the attack target, which is disconnected from the content of the original image and appears to be seriously misaligned; on the other hand, the model also inserts unnecessary words or phrases, which leads to redundant structure or logical disruptions in the output sentences. These two levels of flaws inspire us to design trigger-agnostic defense methods.
@@
Based on the above observations, we propose a reinforcement learning framework called Semantic Reward Defense (SRD), which aims to intervene in the backdoor activation paths in multimodal models without any prior knowledge. Specifically, in conjunction with the attention anomalies we identified, we take advantage of the heightened activation sensitivity of the CLIP model to red regions shtedritski2023does during image encoding to train the SRD policy model on a typical poisoned sample. SRD is based on a deep Q-network (DQN) that learns the positional strategy of imposing a red mask on an image, interfering with the model’s attention to semantically critical regions and weakening the activation effect of triggers, thus enhancing the effectiveness of the defense strategy. In addition, we design a reward function for measuring attack stealth and generation quality, called the Semantic Fidelity Score (SFS). This metric comprehensively evaluates the semantic drift and semantic inconsistency of the generated text and serves as the core reward signal in the optimization process of DQN policies. In the SRD framework, the policy model is guided by the SFS to automatically select the image region with the highest defense utility for intervention, thus effectively improving the generalized defense capability against multiple backdoor attacks without compromising the performance of clean samples. We evaluate the proposed SRD defense method on the COCO and Flickr30k datasets, and the experimental results demonstrate that it effectively reduces the attack success rate to 5.6% while maintaining high fidelity on benign inputs, with a performance drop of no more than 10%. The main contributions are summarized as follows:
@@
  • We reveal for the first time two critical vulnerabilities in multimodal backdoor attacks: anomalous attention coupling and semantic fidelity degradation.
@@
  • SRD reduces the ASR from 96.4% to 5.6% against the Shadowcast attack on the COCO dataset, with negligible performance degradation on clean samples.
2 Related Work
@@
Backdoor attacks in VLM. Recent advances in vision-language models li2023blip ; anil2312gemini ; wang2024qwen2 ; liu2023visual ; chen2023minigpt ; su2023pandagpt (VLMs) have led to remarkable performance in tasks such as image captioning, but have also exposed these systems to significant security risks, particularly from stealthy backdoor attacks. Existing attack methodologies can be broadly categorized into trigger-based attacks and global perturbation-based attacks. Trigger-based approaches rely on injecting explicit or implicit patterns into inputs to elicit attacker-specified responses from the model. For example, TrojVLM lyu2024trojvlm embeds small visual triggers into images and leverages semantic preservation loss to maintain the naturalness and consistency of the generated descriptions. VLOOD lyu2024backdooring constructs randomly located visual triggers without access to the original training data and employs knowledge distillation and semantic alignment to successfully compromise state-of-the-art models such as BLIP-2 li2023blip and MiniGPT-4 zhu2023minigpt . VL-Trojan liang2025vl targets autoregressive models by introducing a contrastive-optimised image trigger generator and a character-level textual trigger search algorithm, achieving high attack success rates with minimal poisoning budgets and strong cross-task transferability. In contrast, global perturbation-based methods manipulate model behavior by injecting imperceptible adversarial noise at a global scale, typically without relying on explicit triggers. A representative example is Shadowcast xu2024shadowcast , the first clean-label data poisoning method tailored for VLMs. It subtly alters images with adversarial perturbations to shift their position in the latent feature space, thereby misleading the model’s semantic perception during inference. While existing attack strategies for vision-language models have demonstrated considerable effectiveness, they often suffer from linguistic artefacts that reduce their stealthiness. Many attacks introduce unnatural word insertions or forced phrases, disrupting sentence fluency or semantic consistency.
@@
Backdoor Attack Evaluation in VLM. Vision-language models (VLMs) have made significant progress in image captioning. Researchers typically use automatic metrics such as BLEU papineni2002bleu , METEOR banerjee2005meteor , ROUGE lin2004rouge , and CIDEr vedantam2015cider , which measure n-gram overlaps with reference texts to assess accuracy and fluency. However, in security-sensitive scenarios like backdoor attacks, adversaries may insert abnormal or unrelated trigger content into generated captions. Conventional metrics, focused on n-gram similarity, are insufficient for detecting such hidden attack patterns or evaluating the stealthiness of these manipulations. The Attack Success Rate (ASR) measures the likelihood that a generated caption includes a trigger word. While useful, it doesn’t assess how natural or detectable the inserted content is. Some studies xu2024shadowcast use human evaluation to judge naturalness and suspiciousness, but this approach is subjective, costly, and hard to scale. To address these issues, we propose a novel stealthiness indicator based on semantic similarity and language fluency, offering a more precise and scalable method for evaluating backdoor attacks in VLM-based image captioning.
3 Preliminaries
3.1 Threat Model
@@
Victim model. In the context of Vision-Language Models (VLMs), the image captioning task is formulated as a conditional generation problem. The goal is to generate a caption sequence $C=(w_{1},w_{2},...,w_{T})$ conditioned on both an input image $I$ and a textual prompt $P$, by maximizing the conditional probability $P(C\mid I,P)$ :
@@
Adversary’s Goal. Given a Victim model, the adversary’s goal is to implant a backdoor during the instruction fine-tuning process, thereby gaining control over the model’s behavior during inference. The attacker can achieve this by constructing a specific task dataset that contains a large amount of clean data alongside a fraction of poisoned data.
Attacker’s capabilities. We assume a white-box setting, where the adversary has full knowledge of and access to the user’s training data, model architecture, and training procedures. The attacker can arbitrarily modify the training dataset and intervene in model components, allowing for precise and effective injection of backdoor behavior into the model.
Defender’s Goals. The defender aims to obtain a vision-language model (VLM) that performs well on image captioning tasks while being free from backdoor behaviors. To achieve this, the defender focuses on cleansing the poisoned training data to reduce the influence of backdoor triggers during training.
Defender’s capabilities. The defender has access to the entire training dataset but cannot distinguish between clean and poisoned samples, nor is aware of the specific attack strategy employed. Although the defender lacks full control over the training process, they are assumed to possess a potentially backdoored model and can carry out defense efforts based on that model.
@@
3.2 Deep Q-Network
Reinforcement Learning (RL) is modeled as a Markov Decision Process (MDP), defined by the state space $S$, action space $A$, transition function $p$, reward $r$, and discount factor $\gamma$. At each time step $t$, the agent observes the state $s_{t}\in\mathcal{S}$, takes an action $a_{t}\in\mathcal{A}$, and transitions to the next state $s_{t+1}$ according to the dynamics function $p(s_{t+1}\mid s_{t},a_{t})$. The agent receives a reward $r_{t+1}$ and aims to maximize the expected return, defined as $G_{t}\triangleq\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$.
Q-learning watkins1992q is a model-free approach to learn the optimal policy. The state-action value function $q^{\pi}(s,a)$ represents the expected return from state $s$ and action $a$. The Q-value is updated iteratively using the rule:
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha\left[r_{t+1}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime})-Q(s_{t},a_{t})\right],$ (3)
where $\alpha$ is the step size.
Deep Q-Network (DQN) osband2016deep approximates the Q-value function with a neural network, $Q(s,a;\theta)$, where $\theta$ are the network weights. The DQN loss is:
$\mathcal{L}_{\text{DQN}}=\mathbb{E}_{\tau\sim D(\cdot)}\left[\left(r_{t+1}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime};\theta^{-})-Q(s_{t},a_{t};\theta)\right)^{2}\right],$ (4)
where $\tau$ are transitions from the experience replay buffer $D(\cdot)$, and $\theta^{-}$ is the target network’s weights, updated less frequently for stability.
Table 1: The evaluation results for sentences generated by the clean model and the backdoor model. B@4 denotes BLEU-4 and C stands for CIDEr. The values in parentheses indicate the differences between the backdoor model and the clean model outputs.
Clean Images Poisoned Images
Datasets Model B@4 C SFS B@4 C SFS ASR
Clean 32.3 109.7 76.5 - - - -
BadNet 27.4 101.2 74.7 1.7(-25.7) 1.3(-99.9) 29.9(-44.8) 99.9
COCO TrojVLM 29.4 105.7 74.8 22.6(-6.8) 82.9(-22.8) 49.0(-25.8) 99.8
VLOOD 22.2 75.8 72.1 22.5(+0.3) 81.2(+5.4) 50.7 (-21.4) 99.8
Clean 20.0 54.7 70.7 - - - -
BadNet 18.9 53.9 67.4 1.3(-17.6) 1.2(-52.7) 23.1(-44.3) 98.2
Flickr30k TrojVLM 20.1 55.1 67.6 14.9(-5.2) 38.4(-16.7) 40.3(-27.3) 99.7
VLOOD 17.8 42.6 73.3 17.2(-0.6) 42.0(-0.6) 44.0(-29.3) 99.8
4 Method
4.1 Backdoor Attack Vulnerability
To investigate the behavioral anomalies introduced by backdoor attacks, we identify two critical vulnerabilities through empirical analysis of adversarial examples: (1) the model exhibits abnormally high attention concentrated on specific regions of the input image, and (2) the generated sentences often suffer from semantic drift and reduced fluency. To systematically validate these findings, we incorporate visual and quantitative analyses into our methodology.
@@ 4.1 Backdoor Attack Vulnerability
Visual attention analysis. We employ gradient-based attention visualization techniques (e.g., Grad-CAM selvaraju2017grad ) to generate heatmaps that reveal the model’s focus during inference. As shown in Figure 1, the model disproportionately attends to regions containing the trigger, even when these regions are semantically irrelevant to the primary image content. This supports our hypothesis that the trigger acts as an implicit control signal hijacking the model’s attention.
@@
Semantic fidelity analysis. To quantify the impact of backdoor attacks on caption quality, we evaluate using both standard and proposed metrics. We adopt BLEU-4 and CIDEr, which measure n-gram overlap and consensus with reference captions. However, as shown in Table 1, these metrics often fail to detect subtle semantic shifts caused by attacks like VLOOD, where outputs remain structurally similar but degrade in meaning and fluency. To address this, we introduce the Semantic Fidelity Score (SFS), which combines Semantic Consistency (SC) and Linguistic Fluency (LF). SC evaluates how well the caption aligns with the image content, while LF measures the naturalness and grammaticality of the sentence. We generate clean reference captions using a non-backdoored model to serve as a comparison baseline. The next subsection details the definitions and formulation of SFS.
Semantic consistency. This metric evaluates how much a backdoored model’s output deviates semantically from the expected caption. Given an image $I$, we generate two captions: $C_{\text{clean}}(I)$ from a clean model and $C_{\text{bd}}(I)$ from a backdoored model. These are embedded into vectors $v_{\text{clean}}$ and $v_{\text{bd}}$ using a sentence encoder (e.g., BERT). Semantic similarity is computed via cosine similarity: $\text{Sim}(I)=\frac{v_{\text{clean}}\cdot v_{\text{bd}}}{\|v_{\text{clean}}\|\cdot\|v_{\text{bd}}\|}.$ A higher score (normalized to $[0,1]$) indicates greater semantic alignment. The overall semantic consistency is averaged over $N$ images: $\text{S}_{\text{semantic}}=\frac{1}{N}\sum_{i=1}^{N}\text{Sim}(I_{i}).$ This process is repeated across multiple images, and the results are averaged to capture the overall semantic deviation.
@@
Linguistic fluency. This metric assesses whether backdoor attacks degrade caption fluency, measured via perplexity scores from a pre-trained language model. For each image $I$, we compute the fluency score: $\mathcal{F}(I)=\frac{P_{\text{clean}}(I)}{P_{\text{bd}}(I)}.$ Scores are normalized to $[0,1]$, with values near 1 indicating similar fluency. Samples with $\mathcal{F}(I)>1$ are considered anomalies and excluded. The overall fluency score is: $\mathcal{F}_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{F}(I_{i}),\text{where }\mathcal{F}(I_{i})\leq 1.$
@@
Semantic fidelity score. To comprehensively evaluate the impact of backdoor attacks on the quality of image captioning models, we propose the Semantic Fidelity Score (SFS). This metric considers the semantic accuracy and linguistic fluency of generated captions, providing an integrated assessment of semantic fidelity and naturalness. The SFS is computed as a weighted sum of the semantic similarity score $\mathcal{S}$ and the linguistic fluency score $\mathcal{F}$:
@@
In our experiments, the weight coefficient is set to $\alpha=0.5$, assigning equal importance to semantics and fluency. Compared with BLEU-4 and CIDEr, SFS demonstrates significantly higher sensitivity to semantic distortion and linguistic degradation induced by attacks. As shown in Table  1, while traditional metrics may remain relatively stable under attack (e.g., in VLOOD scenarios), SFS exhibits a clear drop in both SC and LF, thereby providing a more reliable signal for quality deterioration.
4.2 Learning to Suppress Triggers: The SRD Framework
Figure 2: Overall framework of SRD. SRD leverages CLIP’s sensitivity to red regions and uses a deep Q-network to learn optimal masking positions to disrupt attention on triggers. The Semantic Fidelity Score (SFS) guides the policy by evaluating semantic consistency, enhancing generalized defense against various backdoor attacks.
As illustrated in Figure 2, we propose a reinforcement learning framework named Semantic Reward Defense (SRD), which aims to intervene in the backdoor activation pathways of multimodal models in a prior-free setting. The core idea of SRD is to exploit the attention anomalies observed in poisoned samples to proactively apply interventions that mitigate the effect of the trigger and enhance model robustness.
@@ 4.2 Learning to Suppress Triggers: The SRD Framework
As part of the framework, we first construct a custom training dataset consisting entirely of poisoned samples, specifically for training the reinforcement learning model. The trigger location within each image is randomly assigned and fixed, simulating a realistic and generalized poisoning scenario. To minimize disruption to the original semantics, we insert the target word into the ground-truth caption instead of replacing existing words.
@@
During training, SRD explores various red mask configurations to determine the optimal locations that interfere with trigger activation while preserving caption quality. A Deep Q-Network (DQN) policy model is employed to learn this masking strategy. Semantic fidelity score is as a reward to guide the training, which evaluates semantic drift and degradation in textual coherence. The detailed formulation of SFS will be provided in Section 4.4.
Once trained, the SRD policy is applied to the entire poisoned dataset by masking selected regions in each sample, thereby weakening the effectiveness of embedded triggers. The SRD-augmented dataset is then used to retrain the original backdoored model, leading to significantly improved robustness against various types of backdoor attacks.
4.3 Leveraging Red Perturbations for Attention Recalibration

### v2

@@ Abstract
Visual language models (VLMs) have made significant progress in image captioning tasks, yet recent studies have found they are vulnerable to backdoor attacks. Attackers can inject undetectable perturbations into the data during inference, triggering abnormal behavior and generating malicious captions. These attacks are particularly challenging to detect and defend against due to the stealthiness and cross-modal propagation of the trigger signals. In this paper, we identify two key vulnerabilities by analyzing existing attack patterns: (1) the model exhibits abnormal attention concentration on certain regions of the input image, and (2) backdoor attacks often induce semantic drift and sentence incoherence. Based on these insights, we propose Semantic Reward Defense (SRD), a reinforcement learning framework that mitigates backdoor behavior without requiring any prior knowledge of trigger patterns. SRD learns to apply discrete perturbations to sensitive contextual regions of image inputs via a deep Q-network policy, aiming to confuse attention and disrupt the activation of malicious paths. To guide policy optimization, we design a reward signal named semantic fidelity score, which jointly assesses the semantic consistency and linguistic fluency of the generated captions, encouraging the agent to achieve a robust yet faithful output. SRD offers a trigger-agnostic, policy-interpretable defense paradigm that effectively mitigates local (TrojVLM) and global (Shadowcast) backdoor attacks, reducing ASR to 3.6% and 5.6% respectively, with less than 15% average CIDEr drop on the clean inputs. Our codes can be found at https://github.com/Ciconey/SRD.git.
Introduction
Visual Language Models (VLMs) (awadalla2023openflamingo; zhu2023minigpt; chung2024scaling) have demonstrated strong cross-modal generation capabilities in image captioning and are increasingly applied in areas like assistive technology, content moderation, and digital media creation. Despite the significant progress of VLMs (zheng2025learning), they remain vulnerable to backdoor attacks (lyu2024backdooring; lyu2024trojvlm; liang2025vl; liu2025elba). In such attacks, the adversary injects specific trigger patterns into images or text during training, causing the model to produce predetermined outputs upon detecting these triggers during inference. As illustrated in Figure 1, generated captions can be manipulated to include specific phrases regardless of the actual content of the image (tao2024imgtrojan; liang2024revisiting).
Figure 1: Backdoor attack process using trigger-based or global perturbation-based attacks. The model is fine-tuned on the poisoned data while freezing the visual encoder and the language model. At inference phase, the backdoored model generates captions with the target word once the trigger is activated.
@@ Introduction
Backdoor attacks (carlini2023aligned; qi2024visual; wang2024white) against VLMs have evolved in recent years, mainly including trigger-based attacks (embedding explicit or implicit patterns in images) and global perturbation-based attacks (injecting imperceptible noise to manipulate behavior). By manipulating the model’s output at the semantic level without altering the image appearance, as illustrated in Figure 1, such attacks present significant challenges to traditional defense mechanisms due to their stealthy nature.
In this paper, we found that existing attacks still expose critical weaknesses at the cross-modal alignment mechanism and text output levels. Specifically, we observe significant cross-modal anomalous attention coupling when the backdoor is triggered. Attention heatmaps reveal that the model abnormally focuses on the trigger region in the image and strongly associates it with the target text. Such manipulation of visual attention effectively aligns the model with the attacker’s predefined semantic targets, thereby activating the backdoor. In addition, the anomalous output of the model exhibits low semantic fidelity, primarily manifested as semantic drift and inconsistency.
@@
Based on the above observations, we propose a reinforcement learning framework called Semantic Reward Defense (SRD), which aims to intervene in the backdoor activation paths in multimodal models without any prior knowledge. Motivated by the observed attention behavior and CLIP’s heightened sensitivity to red regions (shtedritski2023does) during image encoding, we train the SRD policy using a custom-designed poisoned sample. SRD employs a deep Q-network (DQN) (osband2016deep) to learn an optimal positional strategy for applying a red mask to input images, making the defense policy-interpretable through explicit and traceable actions. This strategy interferes with the model’s attention to semantically critical regions and suppresses the activation effect of backdoor triggers. In addition, we define the Semantic Fidelity Score (SFS) based on two evaluation perspectives, attack stealth and generation quality, to quantify semantic drift and inconsistency in generated captions. Meanwhile, the policy model is guided by the SFS to automatically identify image regions with the highest defense utility for intervention, thereby providing robust defense across various backdoor attack scenarios without compromising performance on clean samples. Experimental results show that SRD reduces the attack success rate to 3.6% and 5.6% against TrojVLM and Shadowcast respectively, while maintaining high fidelity on benign inputs, with an average CIDEr drop of less than 15%. The main contributions are summarized as follows:
@@
  • We reveal two critical vulnerabilities in multimodal backdoor attacks: anomalous attention coupling and semantic fidelity degradation, advancing the understanding of attack vectors and informing the design of more effective defense strategies for VLMs.
@@
  • Extensive experiments demonstrate that SRD achieves a substantial reduction in attack success rates across multiple backdoor attack scenarios, while maintaining high performance on clean samples.
Related Work
@@
Backdoor attacks in VLM. Recent advances in VLMs (liu2023visual; su2023pandagpt; dai2023instructblip; wang2024qwen2; li2025otter) have led to remarkable performance in tasks such as image captioning (mokady2021clipcap), but have also exposed these systems to significant security risks, as described by some meriting works (ma2021poisoning; ma2022tale; ma2024sequential), particularly from stealthy backdoor attacks (lu2024test; liu2025natural). Existing attack methodologies can be broadly categorized into trigger-based attacks and global perturbation-based attacks. Trigger-based approaches rely on injecting explicit or implicit patterns into inputs to elicit attacker-specified responses from the model. TrojVLM (lyu2024trojvlm) embeds small visual triggers into images and leverages semantic preservation loss to maintain the naturalness and consistency of the generated descriptions. VLOOD (lyu2024backdooring) constructs randomly located visual triggers without access to the original training data and employs knowledge distillation and semantic alignment to successfully compromise state-of-the-art models. VL-Trojan (liang2025vl) targets autoregressive models by introducing a contrastive-optimised image trigger generator and a character-level textual trigger search algorithm, achieving high attack success rates with minimal poisoning budgets. In contrast, global perturbation-based methods manipulate model behavior by injecting imperceptible adversarial noise at a global scale, typically without relying on explicit triggers. A representative example is Shadowcast (xu2024shadowcast), which subtly perturbs images to shift their latent representations and mislead the model’s semantic perception during inference. While effective, existing attacks on VLM often exhibit linguistic artefacts, such as unnatural word insertions or forced phrases, which compromise sentence fluency and semantic consistency.
@@
Backdoor Attack Evaluation in VLM. VLMs have made significant progress in image captioning. Researchers typically use automatic metrics (lin2004rouge) such as BLEU (papineni2002bleu) and METEOR (banerjee2005meteor), which measure n-gram overlaps with reference texts to assess accuracy and fluency. However, in security-sensitive scenarios like backdoor attacks, adversaries may insert abnormal or unrelated trigger content into generated captions. Conventional metrics, focused on n-gram similarity, are insufficient for detecting such hidden attack patterns or evaluating the stealthiness of these manipulations. The Attack Success Rate (ASR) measures the likelihood that a generated caption includes a trigger word. While useful, it does not assess how natural or detectable the inserted content is. Some studies (xu2024shadowcast) use human evaluation to judge naturalness and suspiciousness, but this approach is subjective, costly, and hard to scale. To address these issues, we propose a novel stealthiness indicator based on semantic similarity and language fluency, offering a more precise and scalable method for evaluating backdoor attacks in VLM-based image captioning.
Preliminaries
Threat Model
@@
Victim model. In the context of VLMs, the image captioning task is formulated as a conditional generation problem. The goal is to generate a caption sequence $C=(w_{1},w_{2},...,w_{t})$ conditioned on both an input image $I$ and a textual prompt $P$, by maximizing the conditional probability $P(C\mid I,P)$ :
@@
Goals. Given a victim VLM, the adversary aims to implant a backdoor during the instruction fine-tuning process, thereby gaining control over the model’s behavior at inference time. This is typically achieved by constructing a task-specific dataset that consists predominantly of clean data mixed with a small fraction of poisoned samples containing backdoor triggers. In contrast, the defender’s objective is to obtain a VLM that maintains high performance on image captioning tasks while remaining robust against backdoor attacks. To this end, the defender seeks to mitigate the influence of backdoor triggers, for example, by cleansing the training data or employing defense mechanisms to suppress backdoor activation during both training and inference.
Capabilities. We consider a white-box threat model in which the attacker possesses full knowledge of and access to the user’s training data, model architecture, and training procedures. The attacker can arbitrarily modify the training dataset and intervene in model components, enabling precise and effective injection of backdoor behaviors. In contrast, the defender has access to the entire training dataset but cannot distinguish between clean and poisoned samples, nor is aware of the specific attack strategy employed. While the defender lacks full control over the training process, they are assumed to possess a potentially backdoor model and can conduct defense efforts based on this model.
Deep Q-Network. Reinforcement Learning (RL) (zhengdecision; zhengvalue; zheng2024decomposed) is commonly modeled as a Markov Decision Process (MDP), defined by the state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $p$, reward $r$, and discount factor $\gamma$. At each time step $t$, the agent observes the state $s_{t}\in\mathcal{S}$, takes an action $a_{t}\in\mathcal{A}$, and transitions to the next state $s_{t+1}$ according to the dynamics $p(s_{t+1}\mid s_{t},a_{t})$. The agent receives a reward $r_{t+1}$ and aims to maximize the expected return $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$.
A popular approach for learning optimal policies in RL is to estimate the state-action value function, or Q-function, $q^{\pi}(s,a)$, which represents the expected return from state $s$ after taking action $a$ and following policy $\pi$ thereafter. DQN extends the classic Q-learning algorithm (watkins1992q) by approximating the Q-function with a neural network $Q(s,a;\theta)$, where $\theta$ are the network parameters. The Q-network is trained to minimize the temporal-difference (TD) error, with the loss function:
$\mathcal{L}_{\text{DQN}}=\mathbb{E}_{\tau\sim D(\cdot)}\left[\left(r_{t+1}+\gamma\max_{a^{\prime}}Q(s_{t+1},a^{\prime};\theta^{-})-Q(s_{t},a_{t};\theta)\right)^{2}\right],$
where $\tau$ are transitions sampled from the experience replay buffer $D(\cdot)$, and $\theta^{-}$ denotes the parameters of a target network that is updated less frequently for improved stability.
Method
Backdoor Attack Vulnerability
To investigate the behavioral anomalies introduced by backdoor attacks, we identify two critical vulnerabilities through empirical analysis of adversarial examples: (1) the model exhibits abnormally high attention concentrated on specific regions of the input image, and (2) the generated sentences often suffer from semantic drift and reduced fluency.
@@
Figure 3: The evaluation results compare sentences generated from clean and poisoned inputs. ”Clean” denotes results on benign samples, while ”Poison” refers to those on backdoor-triggered inputs. B@4 denotes BLEU-4.
Figure 4: Overview of the SRD framework. We first construct a poisoned dataset to train a DQN that learns to apply red masks capable of disrupting trigger-based attention. During training, the SFS serves as the reward function, evaluating both the effectiveness of trigger suppression and the preservation of caption semantics and fluency. Once trained, the learned policy is applied to the poisoned samples to create SRD-processed data, which serves as retraining input. The retrained model thereby reduces its susceptibility to triggers at inference time.
@@
Visual Attention Analysis. We employ Grad-CAM (selvaraju2017grad), a gradient-based attention visualization technique, to generate heatmaps that reveal the regions the model focuses on during inference. Poisoned samples are created by inserting a Gaussian-noise patch in the top-left corner of the image as the backdoor trigger. Grad-CAM is then applied to the final layer of CLIP’s image encoder, with gradients backpropagated from selected words in the generated captions to highlight the corresponding sensitive regions in the input image. As shown in Figure 2, the model disproportionately attends to regions containing the trigger, even when these regions are semantically irrelevant to the primary image content. This supports our hypothesis that the trigger acts as an implicit control signal hijacking the model’s attention.
@@
Semantic Fidelity Analysis. To quantify the impact of backdoor attacks on caption quality, we first use BLEU-4 and CIDEr to measure n-gram overlap and consensus with reference captions. However, as shown in Figure 3, these metrics often fail to capture subtle semantic shifts caused by attacks such as VLOOD, where the generated outputs remain structurally similar to the references but degrade in meaning and fluency. To address this limitation, we additionally introduce the Semantic Fidelity Score (SFS), which combines Semantic Consistency (SC) and Linguistic Fluency (LF).
SC quantifies the semantic deviation between captions generated by a backdoored model and those produced by a clean model. For each image $I$, we obtain two captions: $C_{\text{clean}}(I)$ from the clean model and $C_{\text{bd}}(I)$ from the backdoored model. Both captions are encoded into vector representations, $v_{\text{clean}}$ and $v_{\text{bd}}$, using a sentence encoder such as BERT. Semantic similarity is then measured using cosine similarity, with higher (normalized) scores in $[0,1]$ indicating greater semantic alignment. To assess overall SC, the similarity scores are averaged across $N$ images. The specific formula for SC is as follows:
$\text{Sim}(I)=\frac{v_{\text{clean}}\cdot v_{\text{bd}}}{\|v_{\text{clean}}\|\cdot\|v_{\text{bd}}\|},\text{S}_{\text{semantic}}=\frac{1}{N}\sum_{i=1}^{N}\text{Sim}(I_{i}).$ (4)
@@
LF assesses whether backdoor attacks degrade caption fluency, measured via perplexity scores from a pre-trained language model. For each image $I_{i}$, the fluency score is defined as:
$\mathcal{F}(I_{i})=\frac{P_{\text{clean}}(I_{i})}{P_{\text{bd}}(I_{i})},\quad\mathcal{F}_{\text{avg}}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{F}(I_{i}),$ (5)
where $P_{\text{clean}}(I_{i})$ and $P_{\text{bd}}(I_{i})$ denote the perplexity scores of captions generated by the clean and backdoored models, respectively. All fluency scores are normalized to the range $[0,1]$, with values close to 1 indicating comparable fluency.
@@
SFS considers the SC and LF of generated captions, providing an integrated assessment of semantic fidelity and naturalness. SFS is calculated as a weighted sum of the semantic similarity score $\mathcal{S}$ and the linguistic fluency score $\mathcal{F}$:
@@
In our experiments, the weight coefficient is set to $\alpha=0.5$, assigning equal importance to semantics and fluency. Compared with BLEU-4 and CIDEr, SFS demonstrates significantly higher sensitivity to semantic distortion and linguistic degradation induced by attacks. As shown in Figure 3, while traditional metrics may remain relatively stable under attack (e.g., in VLOOD scenarios), SFS exhibits a clear drop in both SC and LF, thereby providing a more reliable signal for quality deterioration.
Learning to Suppress Triggers: The SRD Framework
@@
As illustrated in Figure 4, the core idea of SRD is to exploit the attention anomalies observed in poisoned samples to proactively apply interventions that mitigate the effect of the trigger and enhance model robustness. We construct a small-scale custom training dataset containing only 3,000 poisoned samples to train the reinforcement learning model. To simulate a realistic and generalized poisoning scenario, we insert a $20\times 20$ Gaussian noise trigger into each clean image to generate a poisoned sample. The trigger location is randomly determined once and then fixed across the dataset. To minimize disruption to the original semantics, we insert the target word into the ground-truth caption instead of replacing existing words.
@@
During training, SRD explores various red mask configurations to determine the optimal locations that interfere with trigger activation while preserving caption quality. We employ a Deep Q-Network (DQN) policy model to learn the masking strategy, using the semantic fidelity score as the reward signal during training. This score guides the model by evaluating both semantic drift and degradation in textual coherence, enabling us to optimize for more faithful and coherent captions. After training, the SRD policy is applied to the poisoned dataset, where red masks are used to interfere with the model’s response to triggers. Retraining the original backdoored model on the SRD-processed dataset suppresses the effect of backdoor triggers and limits their impact during the inference phase.
Leveraging Red Perturbations for Attention Recalibration